# Normalized solutions of quasilinear Schrödinger equations with a general nonlinearity

### 02 Mar 2024 v1

@@ 1.1. Background and motivation
where $z:\mathbb{R}\times\mathbb{R}^{N}\to\mathbb{C}$, $W:\mathbb{R}^{N}\to\mathbb{R}$ is a given potential and $l$, $h$ are real functions. Quasilinear equations of the form (1.2) appear naturally in mathematical physics and have been accepted as models of several physical phenomena corresponding to various types of nonlinear terms $l$. When $l\left(s\right)=s$, that is, $\varphi(u)=\sqrt{1+2u^{2}}$, Equation (1.2) arises in plasma physics for the superfluid film equation (see [13]), while when $l(s)=\sqrt{1+s}$, that is, $\varphi(u)=\sqrt{1+u^{2}/(2+2u^{2})}$, it models the self-channeling of a high-power ultrashort laser in matter (see [15]). Such equations were also showed in plasma physics and fluid mechanics [16], in dissipative quantum mechanics [8], in the theory of Heisenberg ferromagnetism and magnons [14, 24], and in condensed matter theory [22]. For more details, we refer to [17, 18]. In the last decades, quasilinear Schrödinger equations have received a considerable attentions from researchers on the existence of nontrivial solutions. In [4], M. Colin and L. Jeanjean considered the following quasilinear Schrödinger equations
@@
The assumption $(H)$ is used to guarantee the mountain pass geometry of $\tilde{\mathcal{J}}$ on $S_{c}\times\mathbb{R}$, where $\tilde{\mathcal{J}}(u,s)=\mathcal{J}(e^{sN/2}u(e^{s}\cdot))$. Then a bounded Palais-Smale sequence of $\mathcal{J}$ restricted on $S_{c}$ is obtained. Via some compactness arguments, the author shows the existence of normalized solutions to problem (1.6). More recently, L. Jeanjean, J. Zhang and X. Zhong [12] introduced a new and non-variational approach to deal with problem (1.6) in the mass-subcritical, mass-critical and mass-supercritical case in one unified way. In particular, the assumption $(H)$ can be removed. For more related results on problem (1.6), we refer to [12] and references therein. When $\varphi(u)=\sqrt{1+2u^{2}}$, problem (1.1) reads as the following form
@@
Different to the case $\varphi\equiv 1$, $4+4/N$ was proved by M. Colin, L. Jeanjean and M. Squassina [5] to be the mass-critical exponent for problem (1.7). In [11]. L. Jeanjean, T. Luo and Z.-Q. Wang considered problem (1.7) with a mass-subcritical nonlinearity $f(u)=|u|^{p-2}u$, $p\in(2+4/N,4+4/N)$ and prove the existence of two solutions if the prescribed $L^{2}$-norm is large enough. In [29], H. Ye and Y. Yu focused on the existence of normalized solutions of problem (1.7) in the mass-critical case. For the mass-supercritical case, we refer to [19], where H. Li and W. Zou adopted one perturbation argument as in [20] to get the existence of the existence of ground state normalized solutions to problem (1.7). By applying the index theory, infinitely many normalized solutions were also obtained. Moreover, they also established the concentration behavior of ground state solutions in the mass-critical case. In [30], via the dual approach in [4] and genus theory, L. Zhang, Y. Li and Z.-Q. Wang constructed multiple normalized solutions of problem (1.7) with $f(u)=|u|^{p-2}u$ in the mass-subcritical case. When $\varphi(u)=\sqrt{1+u^{2}/(2+2u^{2})}$, equation (1.1) is reduced to the quasilinear equation
@@ 1.2. Assumptions and main results
It is known that $U,V$ are non-degenerated (See [12, Proposition 2.1]). Now, our main result reads as follows.
@@ 2. Functional setting and preliminaries
By Lemma 2.1, for any fixed $\lambda>0$, $g_{\lambda}$ satisfies the Berestycki-Lions conditions. As a consequence of [2, Theorem 1], for any $\lambda>0$, problem (2.6) admits a ground state solution, which is positive and radially symmetric. Denoting
@@ 3. Asymptotic behaviors
As a consequence of [12, Theorem 2.2-(ii)], $\tilde{v}\equiv 0$, which contradicts $\tilde{v}\left(0\right)=1$. ∎
@@ 4. Uniqueness for $\lambda$ small or large
$\bullet$ Step 1. For any $\lambda>0$, problem (2.4) admits a positive and radially symmetric solution $u_{\lambda}$. By Theorem 4.1, there exists $\lambda_{0}>0$ such that, up to translation, $u_{\lambda}$ is the unique positive solution of problem (2.4) and then is of mountain-pass-type (see [12, Definition 2.1]). Furthermore, the map $\lambda\mapsto u_{\lambda},\lambda\in(0,\lambda_{0})$ is continuous. That is, $\left\{(\lambda,u_{\lambda}):\lambda\in(0,\lambda_{0})\right\}$ is a curve in ×$\mathbb{R}\times H_{rad}^{1}(\mathbb{R}^{N})$. $\bullet$ Step 2. Let $\widetilde{\mathcal{S}}\subset\mathcal{S}$ be the connected component of $\mathcal{S}$ containing the curve $\left\{(\lambda,u_{\lambda}):\lambda\in(0,\lambda_{0})\right\}$, where $\mathcal{S}$ is given in Section 2. Denoting by $P_{1}:(0,+\infty)\times H_{\text{rad }}^{1}\left(\mathbb{R}^{N}\right)\rightarrow(0,+\infty)$ the projection onto the $\lambda$-component, we show that $P_{1}(\widetilde{\mathcal{S}})=(0,+\infty)$.
@@
Then $\mathbb{T}_{\lambda}:H_{rad}^{1}(\mathbb{R}^{N})\rightarrow H_{rad}^{1}(\mathbb{R}^{N})$ is completely continuous and $u$ is a radial solution of problem (2.4) if and only if $u$ is a fixed point of $\mathbb{T}_{\lambda}$ in $H_{\text{rad }}^{1}(\mathbb{R}^{N})$. For $\lambda\in(0,\lambda_{0})$, recall the local fixed point index

### 05 Mar 2024 v2

@@ 1.1. Background and motivation
where $z:\mathbb{R}\times\mathbb{R}^{N}\to\mathbb{C}$, $W:\mathbb{R}^{N}\to\mathbb{R}$ is a given potential and $l$, $h$ are real functions. Quasilinear equations of the form (1.2) appear naturally in mathematical physics and have been accepted as models of several physical phenomena corresponding to various types of nonlinear terms $l$. When $l\left(s\right)=s$, that is, $\varphi^{2}(u)=1+2u^{2}$, Equation (1.2) arises in plasma physics for the superfluid film equation (see [14]), while when $l(s)=(1+s)^{1/2}$, that is, $\varphi^{2}(u)=1+\frac{u^{2}}{2\left(1+u^{2}\right)}$, it models the self-channeling of a high-power ultrashort laser in matter (see [16]). Such equations were also showed in plasma physics and fluid mechanics [17], in dissipative quantum mechanics [9], in the theory of Heisenberg ferromagnetism and magnons [15, 25], and in condensed matter theory [23]. For more details, we refer to [18, 19]. In the last decades, quasilinear Schrödinger equations have received a considerable attentions from researchers on the existence of nontrivial solutions. In [4], M. Colin and L. Jeanjean considered the following quasilinear Schrödinger equations
@@
The assumption $(H)$ is used to guarantee the mountain pass geometry of $\tilde{\mathcal{J}}$ on $S_{c}\times\mathbb{R}$, where $\tilde{\mathcal{J}}(u,s)=\mathcal{J}(e^{sN/2}u(e^{s}\cdot))$. Then a bounded Palais-Smale sequence of $\mathcal{J}$ restricted on $S_{c}$ is obtained. Via some compactness arguments, the author shows the existence of normalized solutions to problem (1.6). More recently, L. Jeanjean, J. Zhang and X. Zhong [13] introduced a new and non-variational approach to deal with problem (1.6) in the mass-subcritical, mass-critical and mass-supercritical case in one unified way. In particular, the assumption $(H)$ can be removed. For more related results on problem (1.6), we refer to [13] and references therein. When $\varphi^{2}(u)=1+2u^{2}$, problem (1.1) reads as the following form
@@
Different to the case $\varphi\equiv 1$, $4+4/N$ was proved by M. Colin, L. Jeanjean and M. Squassina [5] to be the mass-critical exponent for problem (1.7). In [12]. L. Jeanjean, T. Luo and Z.-Q. Wang considered problem (1.7) with a mass-subcritical nonlinearity $f(u)=|u|^{p-2}u$, $p\in(2+4/N,4+4/N)$ and prove the existence of two solutions if the prescribed $L^{2}$-norm is large enough. In [30], H. Ye and Y. Yu focused on the existence of normalized solutions of problem (1.7) in the mass-critical case. For the mass-supercritical case, we refer to [20], where H. Li and W. Zou adopted one perturbation argument as in [21] to get the existence of the existence of ground state normalized solutions to problem (1.7). By applying the index theory, infinitely many normalized solutions were also obtained. Moreover, they also established the concentration behavior of ground state solutions in the mass-critical case. In [31], via the dual approach in [4] and genus theory, L. Zhang, Y. Li and Z.-Q. Wang constructed multiple normalized solutions of problem (1.7) with $f(u)=|u|^{p-2}u$ in the mass-subcritical case. When $\varphi^{2}(u)=1+\frac{u^{2}}{2\left(1+u^{2}\right)}$, equation (1.1) is reduced to the following quasilinear equation
@@ 1.2. Assumptions and main results
It is known that $U,V$ are non-degenerated (See [13, Proposition 2.1]). Now, our main result reads as follows.
@@ 2. Functional setting and preliminaries
  • (v)
    It suffices to show that $\varphi(t)g_{\lambda}(\phi(t))\leq\phi(t)\frac{\rm{d}}{\rm{d}t}[g_{\lambda}(\phi(t))]$ for $t>0$ small. Obviously,
    $\frac{\rm{d}}{\rm{d}t}[g_{\lambda}(\phi(t))]$ $=\frac{f^{\prime}(t)\varphi(t)-f(t)\varphi^{\prime}(t)}{\varphi^{2}(t)}+\frac{\lambda\varphi^{3}(t)-\lambda\varphi(t)+\lambda t\varphi^{\prime}(t)}{\varphi^{2}(t)}$
    By $(\varphi_{0})$, for any $t\geq 0$, $\phi(t)\leq t\varphi(t)$ and then
    $\phi(t)\varphi^{-2}(t)\left[\varphi^{3}(t)-\varphi(t)+t\varphi^{\prime}(t)\right]\geq\phi(t)\varphi(t)-t.$
    Meanwhile, by $(F_{2})$,
    $\lim_{t\rightarrow 0^{+}}\frac{\phi(t)\varphi(t)f^{\prime}(t)}{f(t)}=\alpha-1>1,$
    which implies that for $t>0$ small,
    $\phi(t)\left[f^{\prime}(t)\varphi(t)-f(t)\varphi^{\prime}(t)\right]\geq\varphi^{2}(t)f(t).$
    Thus, it follows that
    $\phi(t)\frac{\rm{d}}{\rm{d}t}[g_{\lambda}(\phi(t))]\geq f(t)-\lambda t+\lambda\phi(t)\varphi(t)=\varphi(t)g_{\lambda}(\phi(t)),\,t>0\,\,\mbox{small}.$
By Lemma 2.1, for any fixed $\lambda>0$, $g_{\lambda}$ satisfies the Berestycki-Lions conditions. As a consequence of [2, Theorem 1], for any $\lambda>0$, problem (2.6) admits a ground state solution in $H^{1}(\mathbb{R}^{N})$, which is positive and radially symmetric. Denoting
@@
Remark 2.2.
As a consequence of [7, Theorem 2], for any $\lambda>0$, for any positive $C^{2}$ solution $u$ of problem (2.6) with $u(x)\rightarrow 0$ as $|x|\rightarrow\infty$, one can show that $u$ is radially symmetric about some point $x_{0}\in\mathbb{R}^{N}$, that is, $u(x)=u_{0}(|x-x_{0}|)$, where $\frac{\partial u_{0}}{\partial r}<0$ for $r=|x-x_{0}|>0$. In what follows, we assume that $x_{0}=0$. Actually, thanks to the elliptic estimate, $u$ and $|\nabla u|$ decays exponentially at infinity. Set $g_{\lambda}(t)=h_{1}(t)+h_{2}(t),t\geq 0$, where
$h_{1}(t)=\frac{f\left(\phi^{-1}\left(t\right)\right)}{\varphi\left(\phi^{-1}\left(t\right)\right)},\,\,h_{2}(t)=-\lambda\frac{\phi^{-1}\left(t\right)}{\varphi\left(\phi^{-1}\left(t\right)\right)}+\lambda t,$
then by $(F_{2})$, $h_{1}(t)=O(|t|^{\alpha-1})$ as $t\rightarrow 0$ and by $(\varphi_{0})$,
$\lim_{t\rightarrow 0}\frac{h_{2}(t)}{t^{2}}$ $=\lambda\lim_{t\rightarrow 0}\frac{\phi(t)\varphi(t)-t}{t^{2}}$
$=\lambda\lim_{t\rightarrow 0}\frac{\phi(t)[\varphi(t)-1]}{t^{2}}+\lambda\lim_{t\rightarrow 0}\frac{\phi(t)-t}{t^{2}}$
$=\frac{3\lambda}{2}\varphi^{\prime}(0)=0.$
It follows that $g_{\lambda}(t)=O(|t|^{\min\{\alpha-1,2\}})$ as $t\rightarrow 0$. Moreover,
$h_{2}^{\prime}(t)=\lambda\varphi^{-3}(\phi^{-1}(t))\left[\varphi^{3}(\phi^{-1}(t))-\varphi(\phi^{-1}(t))+\phi^{-1}(t)\varphi^{\prime}(\phi^{-1}(t))\right]\geq 0,\,\,t\geq 0,$
which implies that $h_{2}$ is nondecreasing for $t\geq 0$. As for the condition on $h_{1}$ as follows, for some $C>0,p>1$,
$\left|h_{1}(t_{1})-h_{1}(t_{2})\right|\leqslant C|t_{1}-t_{2}|/|\log\min(t_{1},t_{2})|^{p},\quad t_{1},t_{2}\in[0,\max_{x\in\mathbb{R}^{N}}u(x)],$
to prove [7, Theorem 2], it is only used in the proof of [7, Lemma 6.3]. Since $u$ decays exponentially, one can check that [7, Lemma 6.3] still holds in our case and as well as [7, Theorem 2].
@@ 3. Asymptotic behaviors
As a consequence of [13, Theorem 2.2-(ii)], $\tilde{v}\equiv 0$, which contradicts $\tilde{v}\left(0\right)=1$. ∎
@@ 4. Uniqueness for $\lambda$ small or large
$\bullet$ Step 1. For any $\lambda>0$, problem (2.4) admits a positive and radially symmetric solution $u_{\lambda}$. By Theorem 4.1, there exists $\lambda_{0}>0$ such that, up to translation, $u_{\lambda}$ is the unique positive solution of problem (2.4) and then is of mountain-pass-type (see [13, Definition 2.1]). Furthermore, the map $\lambda\mapsto u_{\lambda},\lambda\in(0,\lambda_{0})$ is continuous. That is, $\left\{(\lambda,u_{\lambda}):\lambda\in(0,\lambda_{0})\right\}$ is a curve in ×$\mathbb{R}\times H_{rad}^{1}(\mathbb{R}^{N})$. $\bullet$ Step 2. Let $\widetilde{\mathcal{S}}\subset\mathcal{S}$ be the connected component of $\mathcal{S}$ containing the curve $\left\{(\lambda,u_{\lambda}):\lambda\in(0,\lambda_{0})\right\}$, where $\mathcal{S}$ is given in Section 2. Denoting by $P_{1}:(0,+\infty)\times H_{\text{rad }}^{1}\left(\mathbb{R}^{N}\right)\rightarrow(0,+\infty)$ the projection onto the $\lambda$-component, we show that $P_{1}(\widetilde{\mathcal{S}})=(0,+\infty)$.
@@
Then $\mathbb{T}_{\lambda}:H_{rad}^{1}(\mathbb{R}^{N})\rightarrow H_{rad}^{1}(\mathbb{R}^{N})$ is completely continuous and $u$ is a radial solution of problem (2.4) if and only if $u$ is a fixed point of $\mathbb{T}_{\lambda}$ in $H_{\text{rad }}^{1}(\mathbb{R}^{N})$. For $\lambda\in(0,\lambda_{0})$, thanks to Lemma 2.1-$(v)$, similarly to [13, Lemma 7.2], one can get the local fixed point index